A body of mass 1 kg kept on earth's surface is attracted by earth with a force of 10 N .What will be the force of attraction of the earth experienced bby a satellite of mass 800 kg, which is at an altitude of 3R, where R us the radius of the earth. 

Grade:11

2 Answers

Khimraj
5 years ago
F = mg = mGM/R2
10 = 1*GM/R – (1)
then 
F’ = 800*GM/(4R)2  – (2)
(2)/(1)
F’/10 = 800/16
F’ = 500 N
Hope it clears.
Abhinav Gupta
5 years ago
 
F = mg = mGM/R2
10 = 1*GM/R – (1)
then 
F’ = 800*GM/(4R)2  – (2)
(2)/(1)
F’/10 = 800/16
F’ = 500 N
Hope it clears.well